OpenStudy (anonymous):

Use the following four coordinates to determine which one has a distance of 10 units from point C (1, 2).

OpenStudy (anonymous):

(3, 4) (7, 10) (6, 7) (8, 9) I think the answer is (8,9). But i am not sure.. @Orion1213 Can u help me on this ?

OpenStudy (anonymous):

use distance formula...

OpenStudy (johnweldon1993):

Well lets check using the distance formula \[\large \sqrt{(x_2 - x_1)^2 + (y_2 - y_1)^2}\] \[\large \sqrt{(8 - 1)^2 + (9 - 2)^2}\] \[\large \sqrt{(7)^2 + (7)^2}\] \[\large \sqrt{49 + 49}\] \[\large \sqrt{98}\] Does that = 10?

OpenStudy (anonymous):

yes i think so @johnweldon1993

OpenStudy (anonymous):

all you have to do is plug-in each point in the distance formula until you got 10 units of distance...

OpenStudy (anonymous):

doe sthat mean the answer is (7,10) @Orion1213 @johnweldon1993

OpenStudy (johnweldon1993):

\[\large \sqrt{(8 - 7)^2 + (9 - 10)^2}\] \[\large \sqrt{(1)^2 + (-1)^2}\] \[\large \sqrt{1 + 1}\] \[\large \sqrt{2}\] Are you guessing?All you do is follow that same process each time you want to check a coordinate

OpenStudy (anonymous):

you can do the same process for the two remaining coordinates, one of them is sure to give you 10 units of distance....

OpenStudy (anonymous):

i am not gonna lie, yes i am guessing. I am just so confused with the process and checking the coordinates @johnweldon1993

OpenStudy (anonymous):

and that is the answer...

OpenStudy (anonymous):

what's the answer ? @Orion1213

OpenStudy (anonymous):

the distance formula, as the name implies, is your formula to find the distance between two points given their coordinates (x, y)...

OpenStudy (anonymous):

so (6,7) ?? @Orion1213

OpenStudy (anonymous):

no it's (7,10)...

OpenStudy (anonymous):

oh okay yeah that;s what i said before. Thanks. @Orion1213

OpenStudy (anonymous):

can u please help me with another one? @Orion1213

OpenStudy (anonymous):

Write an equation of a line parallel to y = 1 over 4x − 6 in slope-intercept form that passes through the point (−1, 5). y= 1 over 4x − 1 y= 1 over 4x + 5.25 y= −4x + 5 y= −4x − 1.25

OpenStudy (anonymous):

@johnweldon1993